-10x+4y=-30
5x-2y=15

Answer by jim_thompson5910(35256)

First, multiply both sides of equation two by 2

-10x+4y=-30
5x-2y=15


-10x+4y=-30
2(5x-2y)=2(15)


-10x+4y=-30
10x-4y=30


Now add the equations to get 0 = 0


Since the equation is ALWAYS true, there are an infinite number of solutions.